OpenStudy (anonymous):

2x+3y=9 3y+9x=19 solve for x and y

OpenStudy (anonymous):

2x+3y=9 Ans-y=3-2x/3

OpenStudy (anonymous):

3y+9x=19 Ans-y=19/3-3x

OpenStudy (anonymous):

x=10/7

OpenStudy (anonymous):

y=43/21

OpenStudy (anonymous):

x=2.6 y=1.2